OpenStudy (kitkat1):

Medal for best answer!!!!!! Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely? [1 foot = 12 inches]

OpenStudy (anonymous):

length times width times height will give you the volume of the box

OpenStudy (anonymous):

you should decide what units you want to use i recommend feet

OpenStudy (anonymous):

what is 3 inches in terms of feet?

OpenStudy (frozen15):

1 foot =12 inches 1/12 foot=1inch So 3inches=1/12 *3=1/4foot 9inchees=1/12 *9=3/4foot] So the volume is 6foot*1/4foot*3/4foot=18/16=9/8=1.125 cubic foot

OpenStudy (frozen15):

Then divide that by 0.225.... 1.125 / 0.225 = 5

OpenStudy (frozen15):

medal please!

OpenStudy (kitkat1):

U sure its 5?

OpenStudy (frozen15):

yes or you can go by 1.125 cubic feet :) P.S > thanks for the medal